# Homogeneous Tokenizer Matters: Homogeneous Visual Tokenizer for Remote Sensing Image Understanding

### v1

@@ A. Sparse Task
For sparse prediction, we chose the classification tasks on the NWPU-RESISC45 (20% as trainset) and WHU-RS19 datasets. The image size is 224 $\times$ 224, and the metric used is the top-1 accuracy. The experimental results in Table 1 indicate that HOOK requires only 6 tokens to outperform the standard Patch Embed by 7.08 and 9.48, respectively, on the two datasets. Compared to the baselines, HOOK achieves state-of-the-art performance in terms of both token numbers and classification accuracy.
B. Dense Task
@@
For dense prediction, we selected the semantic segmentation task on the GID5 dataset. The image size is 512 $\times$ 512, and the metric used is the mIoU. The dense adapter module introduced in Section 4.4 addresses the challenge of completing dense tasks with a small number of tokens without additional training parameters. The experimental results in Table 1 demonstrate that HOOK requires only 8 tokens to outperform the standard Patch Embed (11.02). Compared to the baselines, HOOK achieves state-of-the-art performance in terms of both token numbers and segmentation accuracy.
@@ C. Efficiency
In addition to accuracy, the efficiency of the visual tokenizer is also crucial. In our experiments, as shown in Table 2, we compared the changes in overall model efficiency and accuracy when Patch Embed and HOOK were used as visual tokenizers.
Although the structure of HOOK is more complex than that of Patch Embed, HOOK significantly reduces the number of tokens, which greatly improves the efficiency of the backbone network. The results in Table 2 indicate that with a pixel window size of 4 $times$ 4, the efficiency of HOOK is lower than that of Patch Embed. As the pixel window size increases, the model’s efficiency improves, while the accuracy gradually decreases. When the pixel window size is 8 $times$ 8, HOOK demonstrates a distinct advantage in efficiency over Patch Embed in terms of the MACs metric (10.66 vs. 16.86), with only a slight decrease in accuracy compared to the 4x4 version of HOOK (0.17), still outperforming Patch Embed (6.91). This finding suggests that HOOK has distinct advantages over Patch Embed in terms of both accuracy and efficiency.
Table 2: Efficiency of HOOK
Tokenizer
Base token
size(pixels)
MACs(G) Acc.(%)
Patch Embed 16 $\times$ 16 16.86 70.30
HOOK(ours) 32 $\times$ 32 5.87 71.73
16 $\times$ 16 6.80 74.13
8 $\times$ 8 10.66 77.21
4 $\times$ 4 26.10 77.38
@@ 5.3 Visualization of homogeneous tokens
Figure 5: Visualization results of homogeneous visual tokens. The left image shows the original image, while the right image displays the region of the token, with each color representing one region of the token. The images in (a) are obtained from the NWPU-RESISC45 classification dataset, while the images in (b) are obtained from the GID5 semantic segmentation dataset.
@@
Tokenizer Backbone Layers Params(M) NWPU2:8(Acc1) WHU-RS19(Acc1)
@@
In Section 4.3, we mentioned that HOOK introduces a cross-attention mechanism that utilizes $N$ learnable query vectors to aggregate image features into $N$ tokens. We saved the intermediate variable, the attention map, which records the correspondence between the $N$ tokens output by HOOK and the original image features. By visualizing the attention map, we can observe the region of the token.
Figure 5 displays the visualization results, where each color represents one region of the token. The results indicate that each token roughly corresponds to a specific object, achieving the goal of SIRs as basic elements of vision. For more visualization results, please refer to A.
5.4 Stacking up to 12 layers is redundant
The standard ViT model consists of two main modules, Patch Embed and 12 layers of self-attention modules, with the primary computational cost concentrated in the latter. A simple intuition is that if a stronger visual tokenizer replaces Patch Embed, the 12 layers of self-attention modules may become redundant.
To validate this intuition, we conducted tests on classification tasks, and Table 3 displays the experimental results. In Table 3, boldface indicates the best performance, and underlining indicates the second-best performance. The experimental results show that when the number of backbone layers is reduced from 12 layers to 1 layer, HOOK incurs less accuracy loss than models using Patch Embed as the visual tokenizer. Specifically, on the NWPU dataset, Patch Embed incurs a loss of 3.85, while HOOK only incurs a loss of 1.93. On the WHU-RS19 dataset, Patch Embed incurred a loss of 3.26, while HOOK incurred only a loss of 0.65. These experimental results preliminarily confirm our intuition.
HOOK has more parameters than Patch Embed. To further illustrate that the effectiveness of HOOK stems from its homogeneity and adaptability rather than solely from the increase in parameters, we also compared ”Patch Embed + 3 backbone layers” and ”HOOK + 1 backbone layer” in Table 3. The former model has an overall parameter count of 21.89 M, while the latter model has 20.90 M, making their overall number of parameters similar. The results in Table 3 indicate that HOOK outperforms Patch Embed by 7.16 on the NWPU dataset and by 8.71 on the WHU-RS19 dataset, demonstrating that the effectiveness of HOOK does not solely depend on the increase in the number of parameters.
Interestingly, when Patch Embed is deployed as the visual tokenizer and the number of backbone layers is reduced from 12 layers to 3 layers, the model’s accuracy only shows a slight fluctuation, and in some cases, such as in the WHU-RS19 dataset, it even improves. This phenomenon is even more pronounced in the results of the segmentation task in Table 4. Both Patch Embed and HOOK show an increase in accuracy, with Patch Embed showing a more significant improvement, further highlighting the redundancy of the 12-layer backbone network. This conclusion may serve as inspiration for the design of future visual backbone models.
@@ 5.4 Stacking up to 12 layers is redundant
5.5 Ablation
@@
Tokenizer Object Perception Module Params(M) Acc.(%)
@@
When designing the OVM, we referred to the implementation of the SAM model and incorporated its local attention module. In Table 5, we present ablation experiments conducted on the local attention and global attention modules. In Figure 6, we visualized the homogeneous tokens under four different conditions: ”No Attention module”, ”Local Attention only”, ”Global Attention only”, and ”Local & Global Attention”. When either local attention or global attention is lacking, the homogeneity of visual tokens significantly decreases, leading to a decrease in model accuracy. However, when stacking local attention and global attention, HOOK is better able to identify SIRs, and the model’s accuracy reaches its highest point. This finding demonstrates that local and global attention mechanisms play important roles in improving the homogeneity of visual tokens, thereby influencing the performance of HOOK.
@@ 5.6 Limitations of HOOK
(1) Homogeneity: Ideally, a homogeneous visual tokenizer should rely on semantic information in the image when identifying SIRs. However, from a visualization perspective, HOOK does not completely eliminate the influence of local textures and colors and still considers color and texture similarity as one of the criteria for determining SIRs. This feature results in poor visualization performance of HOOK in some complex-textured remote sensing images. Additionally, the limitation imposed by local attention also contributes to the decrease in homogeneity. Figure 7 illustrates three typical scenarios with relatively poor homogeneity: (a) constrained by local attention, (b) overly complex colors, and (c) overly complex textures.
In addition, we did not add any additional guidance or constraints to HOOK. HOOK is only supervised by the downstream task loss function when searching for SIRs. Therefore, we are currently not clear on how adding additional prior assumptions (such as connectivity assumptions) will affect the SIRs and the accuracy of the final model. This is one of the directions we will delve into in the future.
@@
Figure 7: Three typical cases of poor homogeneity: (a) shows that HOOK is limited by the attention range of the local attention module, resulting in poor homogeneity areas appearing in the lower right corner; (b) shows that HOOK is influenced by complex color information and fails to find semantic information in the image; and (c) shows that HOOK is influenced by complex texture information.
(2) Adaptability: HOOK can adjust the number of visual tokens as needed. Experiments have shown that it exhibits good adaptability in tasks involving different granularities and images of different sizes. In theory, the larger is the number of tokens, the finer the original image information that is reflected. However, the empirical experiments in Table 6 show that when we increase the number of tokens in classification tasks, the accuracy of HOOK shows an unstable decrease. This phenomenon indicates that the number of visual tokens may have a more complex nonlinear relationship with the original image and the specific task. This is an issue that we will delve into in the future.
@@ 6.1 The visual tokenizer is the pupil of the machine
We believe that visual tokenizers play an extremely important role in transformer-based visual models and even in multimodal models. If the visual model is the eye of the machine, then the visual tokenizer is the machine’s pupil. Its importance is mainly manifested in the following two aspects:
@@
(2) The visual tokenizer directly affects the granularity and efficiency of the model in understanding images. From the perspective of the original image, visual tokens represent the basic elements of the image. From the model’s perspective, a unimodal visual model uses the meaning of tokens and the relationships among tokens to understand the image, while multimodal models establish relationships among modalities at the token level. Therefore, the quality of the visual tokenizer directly affects the granularity and efficiency of image understanding in visual models and multimodal models.
@@ 6.2 Rethinking HOOK
HOOK belongs to the first routing in Section 3.3: ”Splitting and Merging”. First, under the assumption that ”there are no complete objects in a 4 $\times$ 4 pixel window,” HOOK undergoes extreme splitting of the image. When split to a fine enough level, visual tokens will inevitably satisfy “Same Object Multiple Tokens”. Subsequently, we use a self-attention mechanism to associate seeds belonging to the same SIR and use cross-attention to merge tokens within the same SIR, ultimately achieving “Same Object Same Token”.
@@ 6.3 Future work
In addition to the ”Splitting and Merging” method represented by HOOK, the ”Merging and Splitting” method is equally interesting. One possible implementation is to obtain tokens that can represent all objects in the image and then gradually peel off tokens based on the independence between objects. During this process, we may encounter the entanglement issue between semantically similar objects. Resolving this entanglement is a key challenge that needs to be addressed in this approach.
@@
Furthermore, comparing two alternative routings is another worthwhile direction. For example, both routings will encounter the issue of semantic granularity in the second step. The ”Splitting and Merging” method needs to consider whether tokens that represent the fuselage and wings of an aircraft should be merged into one token during the merging process. On the other hand, the ”Merging and Splitting” method considers whether the token that represents the aircraft should be split into the aircraft and wings. The opposite implementations of the same problem in these two methods may cause fundamental differences in certain application scenarios.
@@
Finally, although we only considered the case of ”Same Object Same Token” as the ideal visual token in the previous analysis, are the scenarios of ”Same Object Multiple Tokens” and ”Same Token Multiple Objects” truly meaningless? For tasks such as human-object interaction detection[57, 58] which are aimed at gaining a deeper understanding of interaction relationships, could using more tokens to describe the edges of objects in the case of ”Same Object Multiple Tokens” be more suitable? We will further explore these interesting questions in future research.
7 Conclusion
@@
Multimodal large-scale language models have brought a revolutionary paradigm shift to the field of remote sensing image understanding. The visual tokenizer, as an important and fundamental component, has long been overlooked or even misunderstood. Starting from the essence of the tokenizer, we propose two fundamental properties that an ideal visual tokenizer should possess: (1) Homogeneity: semantic independent regions (SIR) as the basic elements of vision; (2) Adaptability: the number of tokens can be arbitrarily adjusted to support images of any size and tasks of any granularity. To construct a visual tokenizer that satisfies the above two properties, we rigorously define and analyze the binary relationship between tokens and objects and derive two general paths to obtain homogeneous tokens: ”Splitting and Merging” and ”Merging and Splitting”. Based on the former, we propose a simple HOmogeneous visual tOKenizer, HOOK. The experimental results show that HOOK achieves “Same Token Same Object” and outperforms Patch Embed and other baseline methods in both sparse prediction tasks represented by classification and dense prediction tasks represented by segmentation.
@@
This paper emphasizes the importance of visual tokenizers in visual foundation models and multimodal large language models and provide a preliminary theoretical basis for the construction of visual tokenizers. However, the current theoretical analysis and experimental research are still insufficient. For example, the homogeneity that we propose essentially describes the properties of individual tokens, but we have not discussed in detail the emergent behavior exhibited by combinations of multiple tokens. We hope that our initial work can inspire researchers in the field to pay attention to and participate in an in-depth investigation of visual tokens.

### v2

@@ A. Sparse Tasks
For sparse tasks, we choose classification on the NWPU-RESISC45, WHU-RS19, and NaSC-TG2 datasets; as shown in Table 1, HOOK requires only 6 tokens to outperform the standard Patch Embed by margins of 7.08%, 9.48%, and 1.69%, respectively, on these datasets. Compared with the baseline models, HOOK achieves state-of-the-art performance in terms of both the number of tokens used and the classification accuracy.
B. Dense Tasks
@@
For dense tasks, we focus on semantic segmentation on the GID5 and DGLCC datasets. The dense adapter module, introduced in Section 4.4, addresses the challenge of completing dense tasks with a small number of tokens without requiring additional training parameters. As shown in Table 1, HOOK requires only 8 tokens to outperform the standard Patch Embed, with margins of 11.02% and 0.15% on the GID5 and DGLCC datasets, respectively. Compared with the baseline models, HOOK achieves state-of-the-art performance in terms of both the number of tokens used and the segmentation accuracy.
@@ C. Efficiency
In addition to accuracy, the efficiency of the visual tokenizer is crucial. As shown in Table 1, we compare the numbers of parameters and MACs between HOOK and the baseline methods.
Although the structure of HOOK is more complex than that of Patch Embed, HOOK significantly reduces the number of tokens, which greatly improves the efficiency of the backbone network. The results in Table 1 indicate that as the pixel window size increases, the model’s efficiency improves, whereas the accuracy gradually decreases. In the classification task, HOOK-8 demonstrates a distinct advantage in efficiency over Patch Embed in terms of the MACs metric (10.66 vs. 16.86), with only a slight decrease in accuracy compared with HOOK-4 (0.17 on NWPU) and still outperforming Patch Embed (6.91). This trend can be observed across other datasets and suggests that HOOK has distinct advantages over the baseline methods in terms of both accuracy and efficiency.
5.3 Comparison with pretrained Patch Embed
State-of-the-art visual models are typically pretrained on large-scale datasets. Examples include ViT[17] and MTP[66], which are supervised pretrained on labelled data, as well as DINO-MC[62] and SatMAE[69], which are self-supervised pretrained on unlabelled data. In Table 2, we compare the performance of randomly initialized HOOK and pretrained Patch Embed on the NWPU-RESISC45 and WHU-RS19 datasets. Note that we only load the weights of Patch Embed from the pretrained models to fairly compare the effects of different visual tokenizers, while the backbones are randomly initialized. Following the configurations of the original pretrained models, we compare four backbone architectures at different scales: ViT-Small, ViT-Base, ViT-Large, and ViT-Huge. The main differences among these architectures are shown in Table 3.
Table 3: Comparison of different ViT architectures
Arch. Layers Dimension Heads
ViT-Small 12 384 4
ViT-Base 12 768 8
ViT-Large 24 1024 8
ViT-Huge 32 1280 10
Table 2 demonstrates that the accuracy of the randomly initialized HOOK is superior to that of the pretrained Patch Embed across various backbone architectures. Furthermore, owing to HOOK’s significant reduction in the number of visual tokens, its efficiency advantage becomes more pronounced as the backbone size increases. For example, when the ViT-Huge backbone is used, the model’s number of MACs is only 29.41, which is significantly lower than the standard ViT-Huge’s 161.97. In the context of large multimodal models, this characteristic can greatly reduce training and inference costs.
@@ 5.4 Visualization of homogeneous tokens
Figure 5: Visualization results for homogeneous visual tokens. The left image is the original image, whereas the right image displays the regions of the token, with each colour representing one region of the token. The images in (a) are obtained from the NWPU-RESISC45 classification dataset, whereas the images in (b) are obtained from the GID5 semantic segmentation dataset.
Figure 6: Visualization comparison between regions of the token in HOOK and those in the patch-based tokenizers. Visual tokens in HOOK can adaptively adjust according to the shapes of the objects rather than being regular rectangular patches.
In Section 4.3, we mentioned that HOOK introduces a cross-attention mechanism that utilizes $N$ learnable query vectors to aggregate image features into $N$ tokens. We save the intermediate variable, namely, the attention map, which records the correspondence between the $N$ tokens output by HOOK and the original image features. By visualizing the attention map, we can observe the regions of the tokens.
Figure 5 displays the visualization results, where each colour represents one region of each token. The results indicate that each token corresponds to a specific object and that the regions of the token can adaptively adjust according to the shape of the object rather than being a regular rectangular patch. Figure 6 shows a comparison between regions of the token in HOOK and those in Patch Embed, which demonstrates that HOOK can describe the information of objects in the image more clearly with fewer tokens. For more visualization results, please refer to A.
5.5 Stacking up to 12 layers is redundant
The standard ViT model consists of two main modules, Patch Embed and 12 layers of self-attention modules, with the primary computational cost concentrated in the latter. Intuitively, if a stronger visual tokenizer replaces Patch Embed, the 12 layers of self-attention modules may become redundant.
To validate this idea, we conduct tests on classification tasks, and Table 4 displays the experimental results. In Table 4, boldface indicates the best performance, and underlining indicates the second-best performance. The experimental results show that when the number of backbone layers is reduced from 12 layers to 1 layer, HOOK results in less accuracy loss than models that use Patch Embed as the visual tokenizer. Specifically, on the NWPU dataset, Patch Embed incurs a loss of 3.85, whereas HOOK incurs a loss of only 1.93. On the WHU-RS19 dataset, Patch Embed incurs a loss of 3.26, whereas HOOK incurs a loss of only 0.65. These experimental results preliminarily confirm our intuition.
@@ 5.5 Stacking up to 12 layers is redundant
Tokenizer Layers Params(M) NWPU RS19
@@
HOOK has more parameters than Patch Embed does. To further illustrate that the effectiveness of HOOK stems from its homogeneity and adaptability rather than solely from the increase in the number of parameters, we also compare ”Patch Embed + 3 backbone layers” and ”HOOK + 1 backbone layer” in Table 4. The former model has an overall parameter count of 21.89 M, whereas the latter model has 20.90 M, making their overall numbers of parameters similar. The results in Table 4 indicate that HOOK outperforms Patch Embed by 7.16 on the NWPU-RESISC45 dataset and by 8.71 on the WHU-RS19 dataset, demonstrating that the effectiveness of HOOK does not depend solely on the increase in the number of parameters.
Interestingly, when Patch Embed is deployed as the visual tokenizer and the number of backbone layers is reduced from 12 layers to 3 layers, the model’s accuracy only slightly fluctuates, and in some cases, such as on the WHU-RS19 dataset, it even improves. This phenomenon is even more pronounced in the results for the segmentation task in Table 5. Both Patch Embed and HOOK show increases in accuracy, with Patch Embed showing a more significant improvement, further highlighting the redundancy of the 12-layer backbone network. This conclusion may serve as inspiration for the design of future visual backbone models.
5.6 The deeper the tokenizer is, the better?
Table 6: Results of HOOK with different numbers of OPM and OVM layers
Arch. of OPM Layers of OPM Layers of OVM Params (M) MACs (G) Acc. $\Delta$
L+G 2 1 97.23 26.10 77.38 -
L×$\times$2+G×$\times$2 4 1 103.54 45.85 78.40 +1.02
(L+G)$\times$2 4 1 103.54 45.85 78.41 +1.03
L×$\times$3+G×$\times$3 6 1 109.84 65.61 78.25 +0.87
(L+G)$\times$3 6 1 109.84 65.61 77.60 +0.22
L×$\times$4+G×$\times$4 8 1 116.15 85.36 77.91 +0.53
(L+G)$\times$4 8 1 116.15 85.36 77.82 +0.44
L×$\times$3+G×$\times$3 6 1 109.84 65.61 78.25 +0.87
L×$\times$3+G×$\times$3 6 2 122.32 66.51 77.71 +0.33
L×$\times$3+G×$\times$3 6 3 134.78 67.41 77.30 -0.08
L×$\times$4+G×$\times$4 8 1 116.15 85.36 77.91 +0.53
L×$\times$4+G×$\times$4 8 2 128.63 86.26 78.21 +0.83
L×$\times$4+G×$\times$4 8 3 141.08 87.17 77.40 +0.02
Following Section 5.5, a natural question arises: ”Do deeper tokenizers perform better?” In Table 6, we compare the accuracy of HOOK with different depths on the NWPU dataset.
For the OPM, we test two methods of increasing depth. For example, L×$\times$3+G×$\times$3 represents stacking three layers of local attention followed by three layers of global attention, whereas (L+G)$\times$3 represents alternating local attention and global attention three times. For the OVM, we keep the original Perceiver[57] architecture, stacking self-attention and cross-attention on top of cross-attention. For example, a three-layer OVM consists of one layer of cross-attention followed by two layers of alternating self-attention and cross-attention.
The experimental results in Table 6 indicate that while the deeper HOOK shows slightly better accuracy than the standard version does, the improvement diminishes as the number of layers increases. This suggests that the standard version of HOOK is already the optimal configuration for balancing accuracy and efficiency on the current dataset.
@@ 5.6 The deeper the tokenizer is, the better?
Tokenizer Object Perception Module Params (M) Acc.
@@
5.7 Ablation
When designing the OVM, we refer to the implementation of the SAM model and incorporate its local attention module. In Table 7, we present ablation experiments conducted on the local attention and global attention modules. In Figure 7, we visualize the homogeneous tokens under four different conditions: ”no attention module”, ”local attention only”, ”global attention only”, and ”local & global attention”. When either local attention or global attention is lacking, the homogeneity of the visual tokens significantly decreases, leading to a decrease in model accuracy. However, when local attention and global attention are stacked, HOOK is better able to identify SIRs, and the model’s accuracy reaches its highest point. This finding demonstrates that local and global attention mechanisms play important roles in improving the homogeneity of visual tokens, thereby influencing the performance of HOOK.
@@ 5.8 Limitations of HOOK
Figure 8: Three typical cases of poor homogeneity: (a) shows that HOOK is limited by the attention range of the local attention module, resulting in poor-homogeneity areas appearing in the lower-right corner; (b) shows that HOOK is influenced by complex colour information and fails to find semantic information in the image; and (c) shows that HOOK is influenced by complex texture information.
@@
(1) Homogeneity: Ideally, a homogeneous visual tokenizer should rely on semantic information in the image when identifying SIRs. However, from a visualization perspective, HOOK does not completely eliminate the influence of local textures or colours and still considers colour and texture similarity as one of the criteria for determining SIRs. This feature results in poor visualization performance of HOOK on some complex-textured remote sensing images. Additionally, the limitation imposed by local attention also contributes to the decrease in homogeneity. Figure 8 illustrates three typical scenarios with relatively poor homogeneity: (a) constrained by local attention, (b) overly complex colours, and (c) overly complex textures.
In addition, we do not add any additional guidance or constraints to HOOK. HOOK is only supervised by the downstream task loss function when searching for SIRs. Therefore, we are currently not clear on how adding additional prior assumptions (such as connectivity assumptions) affects the SIRs and the accuracy of the final model. This is one of the directions we will delve into in the future.
@@
(2) Adaptability: HOOK can adjust the number of visual tokens as needed. Experiments have shown that it exhibits good adaptability in tasks involving different granularities and images of different sizes. In theory, the larger the number of tokens is, the finer the original image information that is reflected. However, the empirical experiments in Table 8 show that when we increase the number of tokens in classification tasks, the accuracy of HOOK decreases. This phenomenon indicates that the number of visual tokens may have a more complex nonlinear relationship with the original image and the specific task. This is also an issue that we will delve into in the future.
@@ 6.1 The visual tokenizer is the pupil of the machine
We believe that visual tokenizers play an extremely important role in transformer-based visual models and even in multimodal models. If the visual model is the eye of a machine, then the visual tokenizer is the machine’s pupil. Its importance is manifested mainly in the following two aspects:
@@
(2) The visual tokenizer directly affects the granularity and efficiency of the model in understanding images. From the perspective of the original image, visual tokens represent the basic elements of the image. From the model perspective, a unimodal visual model uses the meaning of tokens and the relationships among tokens to understand the image, whereas multimodal models establish relationships among modalities at the token level. Therefore, the quality of the visual tokenizer directly affects the granularity and efficiency of image understanding in visual models and multimodal models.
@@ 6.2 Rethinking HOOK
HOOK corresponds to the first route in Section 3.3: ”splitting and merging”. First, under the assumption that ”there are no complete objects in a 4$\times$4 pixel window,” HOOK undergoes extreme splitting of the image. When split to a fine enough level, visual tokens will inevitably satisfy the ”same object multiple tokens” scenario. We subsequently use a self-attention mechanism to associate seeds belonging to the same SIR and use cross-attention to merge tokens within the same SIR, ultimately achieving the ”same object same token” scenario.
@@ 6.3 Future work
Compared to the ”splitting and merging” method represented by HOOK, the ”merging and splitting” method is equally interesting. One possible implementation is to obtain tokens that can represent all objects in an image and then gradually peel off tokens on the basis of the independence between objects. During this process, we may encounter an entanglement issue between semantically similar objects. Resolving this entanglement is a key challenge that needs to be addressed in this approach.
@@
Furthermore, comparing the two alternative routes, namely, ”splitting and merging” and ”merging and splitting”, is another worthwhile direction. For example, both routes encounter the issue of semantic granularity in the second step. The ”splitting and merging” method needs to consider whether tokens that represent the fuselage and wings of an aircraft should be merged into one token during the merging process. On the other hand, the ”merging and splitting” method considers whether the token that represents the aircraft should be split into the aircraft and wings. The opposite implementations of the same problem in these two methods may cause fundamental differences in certain application scenarios.
@@
Finally, although we only considered the case of ”same object same token” as the ideal visual token in the previous analysis, are the scenarios of ”same object multiple tokens” and ”same token multiple objects” truly meaningless? For tasks such as human–object interaction detection[71, 72], which are aimed at gaining a deeper understanding of interaction relationships, could using more tokens to describe the edges of objects in the case of ”same object multiple tokens” be more suitable? We will further explore these interesting questions in future research.
7 Conclusions
@@
Multimodal large-scale language models have led to a revolutionary paradigm shift in the field of remote sensing image understanding. Visual tokenizers, as important and fundamental components, have long been overlooked or even misunderstood. Starting from the essence of these tokenizers, we propose two fundamental properties that an ideal visual tokenizer should possess: (1) homogeneity: semantically independent regions (SIRs) are the basic elements of vision and (2) adaptability: the number of tokens can be arbitrarily adjusted to support images of any size and tasks of any granularity. To construct a visual tokenizer that satisfies the above two properties, we rigorously define and analyse the binary relationship between tokens and objects and derive two general paths for obtaining homogeneous tokens: ”splitting and merging” and ”merging and splitting”. On the basis of the former, we propose a simple HOmogeneous visual tOKenizer, HOOK. The experimental results show that HOOK achieves the ”same token same object” scenario and outperforms Patch Embed and other baseline methods in both sparse prediction tasks represented by classification and dense prediction tasks represented by segmentation.
@@
This paper emphasizes the importance of visual tokenizers in visual foundation models and multimodal large language models and provides a preliminary theoretical basis for the construction of visual tokenizers. However, the current theoretical analysis and experimental research are still insufficient. For example, the homogeneity that we propose essentially describes the properties of individual tokens, but we have not discussed in detail the emergent behaviour exhibited by combinations of multiple tokens. We hope that our initial work inspires researchers in the field to give attention to and participate in an in-depth investigation of visual tokens.